From GlueXWiki
Jump to: navigation, search

July 28, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126 (Beni)
  5. Other


Participants: Eugene, Bill, Dave, Chris, Beni, Simon, Caleb, and Lubomir.


- Dave (see the construction tracking): Techs are finishing the cathodes for cell#5 and started working on cell#6 cathodes. The components on wire frame #3 were soldered (except few resistors). Today started tension measurements on wire frame #4.

- Dave: Yesterday we started testing the electroplating. First attempt not successful because of wrong polarity. Following the old prescription resulted in Cu deposition that was too brittle. Later Bill studied the procedure (see xlsx file attached) and concluded that the current must be much lower and correspondingly the plating time bigger. The last sample using his procedure was shown at the meeting and looked much better. So, thanks to David's micro-controller one can easily change these parameters and improve the procedure. Eugene: what is the QA for this? Bill: looking with high magnification microscope (need price estimate), measuring the diameter with micrometer, may try using the laser sensor and Caleb will look into that, and finally testing with HV.

- Lubomir studied the leakage currents between the field and sense wires and it turned out it is the new epoxy (Epolite) to be blamed. On an old wire frame short field and sense wires were soldered (11 of each type) on two separate HV boards in the middle of the frame. They were covered with Epolite on one board and with the old epoxy (Epon, used for the prototype) on the other board. After ~24 hours we applied HV and the leakage current between the field and sense wires at the Epolite side was ~40 times higher (~7 uA at 500V) than on the Epon side. We found also that the current depends strongly on the humidity. That's why after flushing wire frame #2 with nitrogen for 3 days the current went down from ~7 uA at 2000V to ~500 nA. Conclusion: we will switch to the old epoxy since there's no time for investigations. However, we will investigate how to "fix" the first four wire frames made already with Epolie, most likely by drying them in a nitrogen box. Eugene: that's why you don't want to change materials that were used before and were working. We switched to Epolite because it is easier to apply it due to the higher viscosity.

- A short term production schedule (see FastTrack linked above) was proposed, that reflects the above findings. The wire frames built so far will be used as spares and assembled in a spare package. Priority now will be to build wire frames for the first package using the old epoxy. At the same time we will build another two cathode sets for the first package. Expect to have the first package ready by middle/end of September. Then after some tests at Blue Crab it will be moved to EEL126. After finishing the first package we will switch to the spare package that will need additional three cathode sets. Till then we will dry the first four wire frames in the nitrogen box, or try other ways of fixing them. Testing the Humi-seal may require measurements with the prototype in 126 with Sr source. The prototype that has a lot of Humi-seal inside the gas volume, was studied for more than a year and no deterioration in the performance was noticed. No radiation effects were studied, but the expected dose at the place of the HV capacitors is about 1kRad for 10 years, much lower than the typical values of about 100kRad-1MRad at which radiation effects in polymers will be noticed.


- Bill expects to have the strong-back for the assembly by the middle of next week (small delay) and it will be needed to assemble the first package. We discussed if we need also third set of hubs and if we have to order now only two gusset rings needed for the second package, or all the rings together. Bill prefers first to see if the rings are rigid enough and then to order all of them.


- Chris: received the rigid flexes populated but not sure yet if these are all. Tested 90 FDC cables out of 140.

Chamber testing

- Beni showed results (page 571) on the first cell efficiency, ~99%, estimated by using external tracking. Modified the gas system to be able to increase the gas flow, but the drift time spectra look the same, i.e. not as before for 90/10% Ar/CO2. Will study the drift time at fixed locations along the wires.

- Beni modified the back-plane on the VXS crate to be able to work with more F1TDCs. Now it will be possible to connect all the wires from one package if we have the pre-amps.


We said thanks to Caleb for his work as a summer student on the wire position measurements.